### v1

@@ 4.3 Shallow water equations
4.4 Discussion
The results of our numerical experiments are reported [^7] In all settings, hyperparameters are chosen to result in macro-architectures with similar parameter counts for the purpose of comparability. The experimental setup, including the choice of hyperparameters, is outlined in Section C in the Appendix. in Table 2. We observe significant performance gains over the baselines in all three problem settings. In particular, we notice that the best performance in the reported relative $L^{2}$-errors is achieved with the inclusion of both global and local convolutions [^8] Similar results are also observed in the setting of zero-shot super-resolution, which are reported in Section C.4., despite an overall reduction in parameter count with respect to their corresponding FNO/SFNO baselines. This is particularly pronounced for the Navier-Stokes and shallow water equations, where errors can accumulate in autoregressive roll-outs. We attribute this to our models’ inductive bias, allowing it to better capture the fine-grained scales and thus achieve better performance.
The Navier-Stokes problem demonstrates the efficacy of combining both differential and local integral kernels with the global convolution of the FNO. Our proposed model with these three components together outperforms models with only two of these three operations. We note that the Darcy flow problem is meant to motivate the need for our proposed differential kernels (and indeed, our best-performing model achieves $87\%$ lower relative $L^{2}$-error than FNO). Since the groundtruth operator in (11) is a differential operator, we expect (and observe) that for high accuracies, very local (i.e., differential) kernels are needed. In particular, we see that the baseline FNO performs poorly, and the local integral kernels do not provide additional benefits. However, we include them in Table 2 for completeness.
@@ 5 Conclusion
In this paper, we have demonstrated a novel framework for local neural operators. We have shown how convolutional layers can be constrained to realize neural operators that approximate differential operators in the continuous limit. Moreover, we have derived convolutions with local integral kernels from the general notion of an integral transform and the related graph neural operator. Finally, we have constructed localized neural operators on the sphere by using discrete-continuous convolutions Ocampo et al. (2022).
@@ Acknowledgements
M. Liu-Schiaffini is grateful for support from the Mellon Mays Undergraduate Fellowship. J. Berner acknowledges support from the Wally Baer and Jeri Weiss Postdoctoral Fellowship. A. Anandkumar is supported in part by Bren endowed chair and by the AI2050 senior fellow program at Schmidt Sciences.
@@ B.1 General Ideas
The following section outlines the basic ideas behind discrete-continuous convolutions as introduced by Ocampo et al. (2022). To generalize the (continuous) convolution (5) to Lie groups and quotient spaces of Lie Groups, we consider the group convolution (see e.g. Cohen & Welling (2016)).
@@
In some cases, signals are not defined on a group but rather on a quotient space $G/H$, where $H$ is a subgroup of $G$. In such cases, a convolution may still be defined by taking $g\in G/H$. For an example, see spherical convolutions Driscoll & Healy (1994); Ocampo et al. (2022).
@@ Appendix C Implementation details
In all three problems, training is conducted by minimizing the squared $L^{2}$-loss until convergence is reached. Our U-Net baseline is adapted from the model and code of the PDE Arena benchmark Gupta & Brandstetter (2022). For the FNOs Li et al. (2020a) and SFNOs Bonev et al. (2023), we use the implementation in the neuraloperator and torchharmonics libraries. Moreover, for all experiments, GELU activation functions and the Adam optimizer are used.
For all three problems, we trained the FNO/SFNO-based models with varying widths and modes, while keeping the overall number of parameters approximately constant. We present the best results for each problem and macro-architecture in Table 2. For the models with local layers, we found that a larger width and fewer modes can often improve performance. We conjecture that the increased embedding dimension confers additional expressivity to the local kernels. This also suggests that local operators are an important inductive bias for these problems. A detailed experimental setup is outlined for all three datasets in the following subsections.
@@
In the 2D Darcy flow setting, we generate our data as described in Section 4.1. We train and test our models and baselines with data discretized onto a $256\times 256$ regular grid. As a baseline, we compare our proposed models with FNO Li et al. (2020a) and the U-Net architecture of Gupta & Brandstetter (2022). We note that this U-Net architecture is not agnostic to the discretization Kovachki et al. (2021), see also Section 3.1. We compare these baselines against our proposed models: the architecture using convolutions with Fourier (i.e., FNO), differential, and integral kernels (Figure 2), as well as architectures using only FNO and differential kernels or only FNO with the proposed integral kernels. Figure 5 compares the predictions of each model.
@@
Training is conducted by minimizing the squared $L^{2}$-loss for 70 epochs on a single NVIDIA P100 GPU, which is sufficient to achieve convergence on all models. We used a step learning rate decay scheduler, starting at $10^{-3}$ and halving every $10$ epochs. The results are shown in Table 2.
@@
For the 2D Navier-Stokes equations, we use the same experimental setup and dataset as Li et al. (2022). The data is discretized on a $128\times 128$ regular grid. The particular form of the Navier-Stokes equation that we consider is a 2D Kolmogorov flow with a Reynolds number of 5000. We compare the same five models as in the Darcy experiment.
@@
C.3 Shallow water Equations
For the shallow water equations, we use the dataset presented by Bonev et al. (2023). To generate samples, initial conditions are drawn from a Gaussian random process at a resolution of $256\times 512$ on an equiangular lat-lon-grid and then advanced in time using a numerical spectral solver. Physical constants such as the sphere’s radius or the Coriolis force are set to match those of Earth. The numerical solver uses $150$ explicit Euler steps to advance the solution $1$ hour in time. The right-hand side is discretized using the spectral basis provided by the Spherical Harmonics. For a detailed description of the dataset, we refer the reader to Bonev et al. (2023) and the corresponding implementation in the torch-harmonics package.

### v2

@@ 4 Experiments
ground truth
prediction
Figure 5: Ground truth and prediction of the horizontal velocity for the flow past a cylinder. The data is represented on an unstructured mesh, which is visualized in gray color. Our proposed architecture can be readily applied to data on unstructured meshes such as this one. For this figure, we learn the residual to the previous time step.
To validate the effectiveness of local operators, we evaluate the architecture on five PDE problems. Figure 4 and Figure 5 show samples on three of these problems. In all five cases, we incorporate our proposed local operators into existing FNO and SFNO architectures to demonstrate that a significant improvement in performance can be achieved by introducing the inductive bias of local convolutions. This section describes the experimental setting for our problems. We provide further experiments, specific implementation details [^5] Code is available in the neuraloperator library at github.com/neuraloperator/neuraloperator and the torch-harmonics library at github.com/NVIDIA/torch-harmonics., and hyperparameters in Appendix C.
4.1 Darcy flow
$-\nabla\cdot(a\nabla u)=f,\quad u|_{\partial D}=0,$ (10)
$u\mapsto-\nabla\cdot(a\nabla u).$ (11)
Following the setup of Hasani & Ward (2024), we take
$a(x)=\begin{bmatrix}x_{1}^{2}&\sin(x_{1}x_{2})\\ x_{1}+x_{2}&x_{2}\end{bmatrix}$
for $x\in D$. We then compare the performance of our proposed models: FNO in parallel with differential kernels, FNO with integral kernels, and FNO with both kernels. We also compare with a baseline FNO and the U-Net architecture from Gupta & Brandstetter (2022). Moreover, we perform zero-shot super-resolution on this problem; these results can be found in Section C.6.
4.2 Navier-Stokes equations
$\nabla\cdot u$ $=0$
on the spatial domain $D=(0,2\pi)^{2}$ equipped with periodic boundary conditions. In the above, $u$ and $p$ denote the velocity and pressure, and Re denotes the Reynolds number of the flow. We want to learn the solution operator mapping initial conditions $u(\cdot,0)=u_{0}$ to the time-evolved solution in vorticity form $w(\cdot,\tau)$ at time $\tau\in(0,\infty)$, where the vorticity is given by
$w=(\nabla\times u)\hat{z}$ (13)
with $\hat{z}$ being a unit vector normal to the plane.
We consider $m=4$ and $\text{Re}=5000$, and we emphasize that learning the solution operator for such a high Reynolds number is very challenging due to the turbulent nature and small-scale features of the flow. We conjecture that the baseline FNO is prone to over-smoothing over the finer scales, consequently leading to a degradation of performance.
4.3 Diffusion-Reaction equation
Moreover, we consider the 2D Diffusion-Reaction equation from Takamoto et al. (2022) which can, for instance, be used for modeling biological pattern formation. In particular, we want to predict the time-evolution of two non-linearly coupled variables, i.e., the activator $u$ and the inhibitor $v$, solving the equation
$\partial_{t}u$ $=c_{u}\Delta u+R_{u}(u,v),$ (14)
$\partial_{t}v$ $=c_{v}\Delta v+R_{v}(u,v),$ (15)
for the spatial domain $D=(-1,1)^{2}$ with no-flow Neumann boundary condition. The reaction functions $R_{u}$ and $R_{v}$ are given by the Fitzhugh-Nagumo equation
$R_{u}(u,v)=u-u^{3}-k-v,\quad R_{v}(u,v)=u-v$ (16)
and we consider $c_{u}=10^{-3}$, $c_{v}=5\cdot 10^{-3}$, and $k=5\cdot 10^{-3}$. The task is to learn the operator mapping the state $(u,v)$ at consecutive time-steps to the time-evolved state and we can compare against the baselines from Takamoto et al. (2022).
Table 2: Results for the 2D diffusion-reaction equation from PDEBench (Takamoto et al., 2022). We present their U-Net baseline results and metrics. In particular, we report the RMSE errors for the low, medium, and high frequency ranges (fRMSE low/med./high), and the (absolute) RMSE, the maximum error, RMSE errors at the boundary (bRMSE), as well as RMSE for the conserved values (cRMSE).
Model # Params Rel. $L^{2}$-error fRMSE low fRMSE med. fRMSE high RMSE max. error bRMSE cRMSE
U-Net $7.8\cdot 10^{5}$ $8.4\cdot 10^{-1}$ $1.7\cdot 10^{-2}$ $8.2\cdot 10^{-4}$ $5.7\cdot 10^{-2}$ $6.1\cdot 10^{-2}$ $1.9\cdot 10^{-1}$ $7.8\cdot 10^{-2}$ $3.9\cdot 10^{-2}$
FNO $9.3\cdot 10^{5}$ $8.3\cdot 10^{-2}$ $6.2\cdot 10^{-4}$ $5.6\cdot 10^{-4}$ $2.4\cdot 10^{-4}$ $5.2\cdot 10^{-3}$ $7.3\cdot 10^{-2}$ $1.5\cdot 10^{-2}$ $1.2\cdot 10^{-3}$
FNO + loc. int (ours) $8.8\cdot 10^{5}$ $6.3\cdot 10^{-2}$ $4.0\cdot 10^{-4}$ $4.6\cdot 10^{-4}$ $1.5\cdot 10^{-4}$ $3.6\cdot 10^{-3}$ $5.0\cdot 10^{-2}$ $1.0\cdot 10^{-2}$ $\mathbf{4.8\cdot 10^{-4}}$
FNO + diff. (ours) $8.8\cdot 10^{5}$ $3.4\cdot 10^{-2}$ $4.4\cdot 10^{-4}$ $1.9\cdot 10^{-4}$ $\mathbf{6.1\cdot 10^{-5}}$ $1.9\cdot 10^{-3}$ $3.5\cdot 10^{-2}$ $4.4\cdot 10^{-3}$ $1.1\cdot 10^{-3}$
FNO + loc. int. + diff. (ours) $8.9\cdot 10^{5}$ $\mathbf{3.1\cdot 10^{-2}}$ $\mathbf{3.2\cdot 10^{-4}}$ $\mathbf{1.8\cdot 10^{-4}}$ $6.2\cdot 10^{-5}$ $\mathbf{1.7\cdot 10^{-3}}$ $\mathbf{3.3\cdot 10^{-2}}$ $\mathbf{4.2\cdot 10^{-3}}$ $7.4\cdot 10^{-4}$
4.4 Shallow water equations
To test the approach on the sphere, we consider the shallow water equations on the rotating sphere. They represent a system of hyperbolic partial differential equations used to model a variety of geophysical flow phenomena, such as atmospheric flows, tsunamis, and storm surges. They can be formulated in terms of the evolution of two state variables $\varphi$ and $u$ (geopotential height and the tangential velocity of the fluid column), governed by the equations
[…]
$F_{ij}=\varphi u_{i}u_{j}+\frac{1}{2}\varphi^{2},$ (18)
and a source term
$f=-2\Omega x\times(\varphi u),$ (19)
which models the Coriolis force due to the rotation of the sphere with angular velocity $\Omega$.
As baselines, we use a planar U-Net architecture, a spherical U-Net, where convolutions are performed using local integral kernels on the sphere, and an SFNO architecture. On the sphere, we only consider the impact of the integral kernel on the SFNO architecture, as it allows for a natural extension to the sphere (see Section 3.4).
4.5 Flow past a cylinder
Due to the formulation of the discrete-continuous convolutions, our approach can be readily generalized to unstructured meshes. To demonstrate this, we consider the Navier-Stokes equations in a two-dimensional channel with a cylindrical structure and a membrane attached to it (see Rahman et al., 2024). Figure 5 depicts a sample from the validation dataset alongside predictions from our model.
Table 3: Results for the cylinder flow problem in a low-data regime with 250 training samples. Baseline results are taken from Rahman et al. (2024) and all architecture including ours perform a direct prediction (no residual prediction) of the velocity and pressure at the next time step.
Model # Parameters MSE error
GINO $6\cdot 10^{7}$ $2.09\cdot 10^{-2}$
DeepONet $6\cdot 10^{6}$ $1.39\cdot 10^{-1}$
GNN $6\cdot 10^{5}$ $5.00\cdot 10^{-3}$
ViT $3\cdot 10^{7}$ $1.19\cdot 10^{-2}$
U-net $3\cdot 10^{7}$ $9.34\cdot 10^{-2}$
FNO + local integral kernel (ours) $10\cdot 10^{6}$ $\mathbf{2.88\cdot 10^{-3}}$
The dataset considers the flow past the cylinder with Reynolds number $\text{Re}=2000$. Rahman et al. (2024) provide several baselines for a low-data regime (250 samples), where the task is to predict the velocity and pressure after a given time step. To deal with the unstructured grid, we adapt the architecture to utilize local integral operators as encoders and decoders to transform the data into a latent representation on an equidistant grid. Then, we can use an FNO with our local integration layers in the latent space.
@@ 4.5 Flow past a cylinder
4.6 Results and discussion
The results of our numerical experiments are reported [^6] In all settings, hyperparameters are chosen to result in macro-architectures with similar parameter counts for the purpose of comparability. The experimental setup, including the choice of hyperparameters, is outlined in Appendix C in the Appendix. in Tables 4, 3 and 2. We observe significant performance gains over the baselines in all five problem settings. In particular, we notice that the best performance in the reported relative $L^{2}$-errors is achieved with the inclusion of both global and local operations, despite an overall reduction in parameter count with respect to their corresponding FNO/SFNO baselines. This is particularly pronounced for the Navier-Stokes and shallow water equations, where errors can accumulate in autoregressive roll-outs. We attribute this to our models’ inductive bias, allowing it to better capture the fine-grained scales and thus achieve better performance. In Table 2, we also show that our local layers achieve lower errors at high frequencies for the diffusion-reaction equation, indicating that local, high-frequency features are correctly captured.
The Navier-Stokes problem demonstrates the efficacy of combining both differential and local integral kernels with the global convolution of the FNO. Our proposed model with these three components together outperforms models with only two of these three operations. While hyperbolic PDEs likely require only local receptive fields (due to finite information propagation (LeVeque, 1992)) and elliptic problems instead require global information, we observe that many practical problems involving a mixture of operators benefit from a hybrid approach such as ours.
We note that the Darcy flow problem is meant to motivate the need for our proposed differential kernels (and indeed, our best-performing model achieves $87\%$ lower relative $L^{2}$-error than FNO). Since the ground truth operator in (11) is a differential operator, we expect (and observe) that for high accuracies, very local (i.e., differential) kernels are needed. In particular, we see that the baseline FNO performs poorly, and the local integral kernels do not provide additional benefits. However, we include them in Table 4 for completeness.
Finally, we perform super-resolution experiments, in which we train the model at a given resolution and then evaluate it at another resolution without fine-tuning. Figure 9 depicts super-resolution results at twice the training resolution for both the Darcy and shallow water equations. The corresponding numerical results are listed in Table 5. We observe that our approach generalizes well across resolutions and outperforms the baselines across all tested resolutions. For a detailed discussion of the super-resolution experiments, we refer the reader to Section C.6.
@@ 5 Conclusion
In this paper, we have demonstrated a novel framework for local neural operators. We have shown how convolutional layers can be constrained to realize neural operators that approximate differential operators in the continuous limit. Moreover, we have derived convolutions with local integral kernels from the general notion of an integral transform and the related graph neural operator. Finally, we have constructed localized neural operators on the sphere by using discrete-continuous convolutions (Ocampo et al., 2022).
@@ Acknowledgements
The authors thank Md Ashiqur Rahman for useful discussions about benchmarks and baselines. M. Liu-Schiaffini is grateful for support from the Mellon Mays Undergraduate Fellowship. J. Berner acknowledges support from the Wally Baer and Jeri Weiss Postdoctoral Fellowship. A. Anandkumar is supported in part by Bren endowed chair and by the AI2050 senior fellow program at Schmidt Sciences.
Impact Statement
The aim of this work is to advance the field of machine learning and scientific computing. While there are many potential societal consequences, none of them are immediate to require specifically being highlight here.
@@ A.1 Theoretical construction
A.2 Empirical evaluation
(a) Convergence
(b) Ground truth
(c) Low-res prediction
(d) Medium-res prediction
(e) High-res prediction
Figure 6: Empirical evaluation of the proposed differential kernel. (a) $L^{2}$ errors at various resolutions and quadratic coefficient scales $c$. (b) True differential operator for $c=1$. (c) Output of the differential kernel at a resolution of $32\times 32$. (d) Output of the differential kernel at a resolution of $64\times 64$. (e) Output of the differential kernel at a resolution of $4096\times 4096$.
In this section, we empirically evaluate the convergence of our proposed differential kernels to a differential operator as the resolution increases. In particular, we consider a randomly-initialized differential kernel satisfying the constraints discussed in Section 3.2 (mean zero and scaled by the resolution $\frac{1}{h}$, where $h$ is the width of each grid cell in an equidistant grid).
To be able to compute the theoretical differential operator in closed form, we consider a simple parabola $v\colon[0,1]^{2}\to\mathbb{R}^{n}$, given by
$x\mapsto\|x\|^{2}\begin{bmatrix}c_{1},\dots,c_{n}\end{bmatrix}^{\top},$ (26)
with coefficients $c_{1},\dots,c_{n}$.
We define a randomly-initialized first-order differential kernel $K=(K_{ij})_{i,j=1}^{3,3}\subset\mathbb{R}^{n}$ (i.e., one differential kernel for each of the $n$ input channels) subject to the constraints described in Section 3.2. Let $b_{j}\in\mathbb{R}^{n}$ denote the direction corresponding to the directional derivative corresponding to the differential kernel for channel $j$, see Section A.1. With the parabola as defined in (26), the output function of the true differential operator corresponding to $K$ is given by
$x\mapsto 2\sum_{i=j}^{n}x\cdot b_{j}.$ (27)
In this simple experiment, we set $n=10$ and generate $c_{1},\dots,c_{n}$ uniformly on the unit interval, multiplied by some scaling factor $c\in\{1,2,4,16\}$. We discretize the parabola in (26) over different resolutions, convolve the kernel $K$ with each of these discretizations, and compute the $L^{2}$ error with respect to the true differential operator. Figure 6 shows the $L^{2}$ error for these values of $c$ and resolutions, as well as visualizations of the true differential operator and the one outputted by the differential kernel at three different resolutions.
From Figure 6, we observe a clear convergence to the differential operator as the resolution increases. For a fixed resolution, convergence is slower for greater magnitudes of the second derivative (i.e., larger $c$). This can be theoretically shown by estimating the remainder of the Taylor expansion in Section A.1.
@@ B.1 General Ideas
The following section outlines the basic ideas behind discrete-continuous convolutions as introduced by (Ocampo et al., 2022). To generalize the (continuous) convolution (5) to Lie groups and quotient spaces of Lie Groups, we consider the group convolution (see e.g. Cohen & Welling (2016)).
@@
In some cases, signals are not defined on a group but rather on a quotient space $G/H$, where $H$ is a subgroup of $G$. In such cases, a convolution may still be defined by taking $g\in G/H$. For an example, see spherical convolutions (Driscoll & Healy, 1994; Ocampo et al., 2022).
@@ Appendix C Implementation details
For Darcy flow, Navier-Stokes, and the shallow water equations, training is conducted by minimizing the squared $L^{2}$-loss until convergence is reached. Our U-Net baseline is adapted from the model and code of the PDE Arena benchmark (Gupta & Brandstetter, 2022). For the FNOs (Li et al., 2020a) and SFNOs (Bonev et al., 2023), we use the implementation in the neuraloperator and torch-harmonics libraries. Moreover, for all experiments, GELU activation functions and the Adam optimizer are used.
For these three problems, we trained the FNO/SFNO-based models with varying widths and modes, while keeping the overall number of parameters approximately constant. We present the best results for each problem and macro-architecture in Table 4. For the models with local layers, we found that a larger embedding dimension and fewer modes can often improve performance. We conjecture that the increased embedding dimension confers additional expressivity to the local kernels. This also suggests that local operators are an important inductive bias for these problems. When supplementing the FNO/SFNO blocks with additional branches, we also observed improved convergence by scaling the initial parameters or the output by $n^{-1/2}$, where $n$ is the number of branches. A detailed experimental setup is outlined for these three datasets in the following subsections. We also present results for the 2D diffusion-reaction equation and for super-resolution experiments on Darcy flow and the shallow water equations.
@@
In the 2D Darcy flow setting, we generate our data as described in Section 4.1. For the input functions, we consider random linear combinations of eigenfunctions of the Laplace operator with zero Dirichlet boundary conditions, i.e.,
$u(x)=\sum_{i,j=1}^{20}\frac{c_{ij}}{\sqrt{(i\pi)^{2}+(j\pi)^{2}}}\sin(i\pi x_{1})\sin(j\pi x_{2}),\quad x\in D,$
with i.i.d. $c_{ij}\sim\mathcal{N}(0,1/(i+j))$. We train and test our models and baselines with data discretized onto a $256\times 256$ regular grid. We use $10000$ training samples and $2000$ testing samples. As a baseline, we compare our proposed models with FNO (Li et al., 2020a) and the U-Net architecture of Gupta & Brandstetter (2022). We note that this U-Net architecture is not agnostic to the discretization (Kovachki et al., 2021), see also Section 3.1. We compare these baselines against our proposed models: the architecture using convolutions with Fourier (i.e., FNO), differential, and integral kernels (Figure 2), as well as architectures using only FNO and differential kernels or only FNO with the proposed integral kernels. Figure 8 compares the predictions of each model.
@@
Training is conducted by minimizing the squared $L^{2}$-loss for 70 epochs on a single NVIDIA P100 GPU, which is sufficient to achieve convergence on all models. We use a step learning rate decay scheduler, starting at $10^{-3}$ and halving every $10$ epochs. The results are shown in Table 4.
@@
For the 2D Navier-Stokes equations (Kolmogorov flows), we use the same experimental setup and dataset as Li et al. (2022a), which sets $m=4$ and $\textrm{Re}=5000$ in (12) and uses a temporal discretization of $\tau=1$ on a $128\times 128$ regular grid. The initial conditions are sampled from a Gaussian measure as described in Li et al. (2022a), and the equation is solved with a pseudo-spectral solver. We compare the same five models as in the Darcy experiment.
@@
C.3 Diffusion-Reaction equation
For the 2D Diffusion-Reaction equation, we use the same experimental setup and dataset as Takamoto et al. (2022). The reference solution is computed using a finite-volume method in space and a fourth-order Runge-Kutta method in time. The dataset consists of $900$ training samples and $100$ validation samples discretized on a $128\times 128$ regular grid with $100$ equidistant time-steps in the interval $[0,5]$ and Gaussian initial conditions. The task is to predict the state of the variables $(u,v)$ at the next time-step from the states at the previous $10$ steps. The errors are measured for the full autoregressive roll-out as in the implementation of PDEBench at github.com/pdebench/PDEBench.
We use the same experimental setup and implementation as Takamoto et al. (2022) and only adapt the learning rate, the number of modes, and the embedding dimension. We then add our proposed layers to (a subset of) the FNO blocks and experiment with both reflective and replicate padding. In particular, we parametrize the local integral kernels by five radial, piecewise linear basis functions and use $3\times 3$ convolutional kernels for the differential kernels.
We train on a single NVIDIA RTX 4090 GPU for 500 epochs with early stopping (using the same criteria as Takamoto et al. (2022)). Moreover, we use an exponentially decaying learning rate, starting at $10^{-4}$ and halving every 100 epochs. We present our best results in Table 2 and refer to Takamoto et al. (2022) for details on the metrics. We compare against the baselines by Takamoto et al. (2022) and note that we also improve their FNO baseline results. However, our local integral and differential kernels still provide a significant improvement with fewer parameters. Specifically, we reduced the modes from $24$ to $16$ and increased the embedding dimension from $20$ to $29$.
C.4 Shallow water Equations
For the shallow water equations, we use the dataset presented by Bonev et al. (2023), which uses a Gaussian random field to generate initial conditions on an equiangular lat-lon-grid on the sphere at a resolution of $256\times 512$ and solves for $\varphi$, $u$ at a lead time of one hour. The target solution is computed using a spectral solver, which takes 24 Euler steps [^8] Dataset and solver are taken from the torch-harmonics package at github.com/NVIDIA/torch-harmonics.. Physical constants such as the sphere’s radius or the Coriolis force are set to match those of Earth. The numerical solver uses $150$ explicit Euler steps to advance the solution $1$ hour in time. The right-hand side is discretized using the spectral basis provided by the Spherical Harmonics. For a detailed description of the dataset, we refer the reader to Bonev et al. (2023) and the corresponding implementation in the torch-harmonics package.